OpenStudy (anonymous):

the fifth term in the expansion of (3x-x^2y)^6 is? the answer -18x^4y^5?

OpenStudy (raden):

actually, it is 6C4(3x)^(6-4)(-x^2y)4 = 6C4 (3x)^2 (-x^2y)4 = 6C4 * 9 x^2 * x^8y^4 = 15 * 9 x^10y^4 = 135 x^10y^4